OpenStudy (anonymous):

To pass a true or false test a student must answer 8 out of 10 correct. What is the probability a student will pass if he quesses on all 10 questions? Find the mean and standard deviation of correct answers. Is it unusual for a student to pass by guessing?

OpenStudy (perl):

need help ?

OpenStudy (anonymous):

OMG U WATCH NARUTO

OpenStudy (anonymous):

haha yea to both

OpenStudy (perl):

we want P( x > = 8) = P( x = 8 ) + P(x = 9) + P(x=10)

OpenStudy (anonymous):

i already got the first question but dont know the other two ):

OpenStudy (perl):

= 10 choose 8 * .5^10 + 10 choose 9 * .5^10 + 10 choose 10 * .5^10

OpenStudy (perl):

for a binomial distribution mean = n*p , so 10*.5 = 5

OpenStudy (perl):

and st. dev is sqrt( n * p * (1-p)) = sqrt( 10 * .5 * .5 )

OpenStudy (perl):

so the mean is 5 answers, the st.dev is 1.5811

OpenStudy (perl):

the probability of passing the test due to guessing is .05468 . I would say it is pretty unlikely

OpenStudy (anonymous):

in other word, u didnt choose the thug lyfe, thug lyfe chose u.

OpenStudy (anonymous):

why is it unusual?

OpenStudy (perl):

because the probability of passing (getting 8 or more right) purely by guessing is so low

OpenStudy (anonymous):

it says to explain why..

OpenStudy (anonymous):

i got a high probability.... i think i made a mistake..

OpenStudy (anonymous):

i understand! ty!

OpenStudy (perl):

go back and do part a) again

OpenStudy (perl):

P( x > = 8) = P( x = 8 ) + P(x = 9) + P(x=10) = 10 choose 8 * .5^10 + 10 choose 9 * .5^10 + 10 choose 10 * .5^10

OpenStudy (anonymous):

i understand lol i did x< =8 instead of x>=8 xD